Hi all, following a column change we decided it was about time to clean the ion source (given that it had been a few months since cleaning, not that we were seeing a significant drop in sensitivity).
Since cleaning the EI source, the emission current, which we are typically seeing tuning at 10-25uA, is now tuning repeatedly at 35uA. Even after a second cleaning of the ion source. The system has also tuned at 35uA with the second filament used during the tune. The abundance/counts of the target ions are also approximately a third of what we typically see. 40,000 cps of the 69 ion, vs a typical abundance of 130,000.
Any ideas are much appreciated!